OpenStudy (anonymous):

solve for x : 3 over x+1= x over 2

OpenStudy (anonymous):

3 x - = - x+1 2 cross multiply 6= x(x+1) 6= x^2 + x x^2+x-6=0 (x+3)(x-2)=0 x= -3 x=2